# supplementary

### v1

@@ 4.1 Experimental Setup
Datasets. Following CoOp (Zhou et al. 2022b) and CoCoOp(Zhou et al. 2022a), experiments under four settings, including cross-dataset generalization, base-to-new generalization, and few-shot classification, are conducted on 11 image classification datasets. Those datasets include ImageNet (Deng et al. 2009) and Caltech101 (Fei-Fei, Fergus, and Perona 2004) for generic object classification, OxfordPets (Parkhi et al. 2012), StanfordCars (Krause et al. 2013), Flowers102 (Nilsback and Zisserman 2008), Food101 (Bossard, Guillaumin, and Van Gool 2014), and FGVCAircraft (Maji et al. 2013) for fine-grained visual categorization, SUN397 (Xiao et al. 2010) for scene recognition, DTD (Cimpoi et al. 2014) for texture classification, EuroSAT (Helber et al. 2019) for satellite image classification, and UCF101 (Soomro, Zamir, and Shah 2012) for action recognition. Besides, following (Zhou et al. 2022a), experiments under the domain generalization setting are conducted on the ImageNet dataset and its variants, including ImageNetV2 (Recht et al. 2019), ImageNet-Sketch (Wang et al. 2019), ImageNet-A (Hendrycks et al. 2021b), and ImageNet-R (Hendrycks et al. 2021a).
@@
Baselines. We evaluate the effectiveness of the proposed ATLaS method by incorporating it into a broad spectrum of baseline methods, including both textual prompt tuning methods (e.g., CoOp (Zhou et al. 2022b), Conditional Context Optimization (CoCoOp) (Zhou et al. 2022a), and Decoupled Prompt Tuning (DePT) (Zhang et al. 2024)) and multi-modal prompt tuning methods (e.g., Independent Vision-Language Prompting (IVLP) (Rasheed et al. 2023) and Multi-modal Prompt Learning (MaPLe) (Khattak et al. 2023a)).
@@
Implementation Details. Our implementation is based on CoOp (Zhou et al. 2022b). For all baseline methods, we maintain the same experimental settings (e.g., training epochs, training schedules, and data augmentation settings) as specified in their original implementations. All experimental results are averaged over three seeds. All experiments are conducted with a ViT-B/16-based CLIP model. The length of prompt vectors is fixed to 4, and they are initialized with the textual template “a photo of a”. For the ATLaS method, $K$, $\theta$ in Eq. (3), $\tau_{c}$ in Eq. (4), and $\alpha$ in Eq. (5) are set to 2, 0.1, 0.05, and 0.1, respectively, by default. More details are provided in Appendix B.
4.2 Experimental Results
In this section, we validate the effectiveness of the proposed ATLaS method that uses the vanilla soft labels defined in Eq. (3) and its two variants, including ATLaS-CSL that uses the CSL to generate soft labels and ATLaS-ISL that uses ISL to generate soft labels, under the cross-dataset generalization, domain generalization, base-to-new generalization, and few-shot classification settings. More details are provided in Appendices C and D.
@@ 4.2 Experimental Results
ImageNet Caltech OxfordPets StanfordCars Flowers Food Aircraft SUN DTD EuroSAT UCF Avg.
@@
Table 1: Performance of various methods under the cross-dataset generalization setting. The number in the parentheses denotes the improvement over the baseline method.
Cross-Dataset Generalization. In the cross-dataset generalization setting that evaluates whether prompts learned from a source domain can generalize to unseen target domains, prompts are trained using 16-shot samples from each of the 1000 classes on ImageNet and then evaluated on other 10 datasets. Results in Table 1 show that CoOp with ATLaS, ATLaS-CSL, and ATLaS-ISL outperform CoOp by 1.09%, 2.10%, and 1.36% in average accuracy, respectively. Unlike ATLaS-ISL which uses instance-class associations from the source domain to generate soft labels, ATLaS-CSL derives the soft labels from inter-class similarities of textual prompts without involving source domain samples. This allows ATLaS-CSL to achieve better cross-dataset generalization capability. Due to the page limit, more results of other baseline methods can be found in Appendix C.1.
Domain Generalization. In the domain generalization setting, prompts are evaluated for the ability to generalize to out-of-distribution data. Specifically, prompts are trained using 16-shot samples from each of the 1000 classes on ImageNet, and then evaluated on four different target domains (i.e., ImageNetV2, ImageNet-Sketch, ImageNet-A, and ImageNet-R). As shown in Table 2, CoOp with ATLaS, ATLaS-CSL, and ATLaS-ISL outperform CoOp by 1.03%, 1.07%, and 0.79% in terms of the average accuracy, respectively. Those results show that the prompts learned by the proposed ATLaS methods are more domain-generalizable. Due to the page limit, more results of other baseline methods can be found in Appendix C.2.
@@
Table 2: Performance of various methods under the domain generalization setting.
@@
Table 3: The average performance of five prompt tuning methods with or without ATLaS, ATLaS-CSL, and ATLaS-ISL on the 11 datasets under the base-to-new generalization setting. ‘tp’ denotes textual prompt tuning, and ‘mmp’ denotes multi-modal prompt tuning. ‘H’ denotes the harmonic mean accuracy.
Base-to-New Generalization. In the base-to-new generalization setting, all classes are divided into disjoint base and new groups, and the prompts are trained on base classes and evaluated on base and new classes. We integrate ATLaS, ATLaS-CSL, and ATLaS-ISL with several baseline methods, including CoOp, CoCoOp, IVLP, MaPLe, and DePT. The experimental settings for all baselines are the same as those in the original implementations. Table 3 shows that the proposed methods consistently enhance the generalization performance of all baselines, specifically achieving better new-class accuracy and harmonic mean accuracy, where full results can be found in Appendix C.3. Those results indicate that ATLaS, ATLaS-CSL, and ATLaS-ISL can be seamlessly integrated with existing prompt tuning methods to enhance the generalization performance of prompts reliably.
Few-shot Classification. We conduct few-shot classification in Appendix C.4, where CoOp with ATLaS, ATLaS-CSL, and ATLaS-ISL all outperform CoOp in terms of the average performance on the 11 datasets.
4.3 Ablation Studies
@@
In this section, we conduct ablation studies to analyze the effectiveness of the alternating mechanism and the proposed soft labels (i.e., CSL and ISL) in the proposed ATLaS method. The experiments are performed using CoOp (Zhou et al. 2022b) as the baseline method under the base-to-new generalization setting.
@@
Table 4 shows the average performance of CoOp equipped with different LS strategies on the 11 datasets. Prompts are trained with 16 instances per base class. Based on the results, we have several observations.
@@
1) CoOp with vanilla LS degrades the performance of CoOp (i.e., -0.45% in the base-class accuracy, -0.91% in the new-class accuracy, -0.75% in the harmonic mean accuracy), suggesting that such a naive combination makes the inconsistency of labels used in the pre-training and fine-tuning phases, weakening generalization performance. In contrast, CoOp with ATLaS, which alternately uses one-hot and soft labels to supervise prompt tuning, outperforms CoOp by 1.96% in the new-class accuracy (i.e., 65.19% vs. 63.23%) and 1.14% in the harmonic mean accuracy (i.e., 72.80% vs. 71.66%), demonstrating that such alternating fine-tuning can enhance prompt generalization.
@@
2) Supervising prompt tuning with CSL and ISL as soft labels improves the new-class accuracy (i.e., +2.5% for CSL, +10.41% for ISL) and harmonic mean accuracy (i.e., +0.64% for CSL, 3.58% for ISL) at the expense of decreasing the base-class accuracy (i.e., -2.36% for CSL, -5.79% for ISL). This occurs because both CSL and ISL leverage the inherent properties of CLIP, allowing them to enhance prompt generalization while preserving the inherent knowledge of CLIP. Building upon this foundation, the alternating mechanism further improves their performance, that is, compared with CoOp, ATLaS-CSL and ATLaS-ISL improve the new-class accuracy by +4.82% and +10.39% respectively, and harmonic mean accuracy by +2.73% and +5.13%. Those results indicate that both the alternating mechanism and soft labels can enhance prompt generalization.
@@
3) Alternating optimization outperforms joint optimization when using both one-hot labels and soft labels to train prompts. That is, CoOp with ATLaS outperforms CoOp with ‘vanilla LS + y’ in terms of the harmonic mean accuracy (i.e., 72.80% v.s. 71.68%), further demonstrating the effectiveness of the proposed alternating mechanism.
Figure 3 shows the test accuracy of CoOp and CoOp with vanilla LS, ATLaS, CSL, ATLaS-CSL, ISL, and ATLaS-ISL on new classes of the SUN397 dataset over the training process. The test accuracy of CoOp first increases and then drops sharply. Integrating CoOp with vanilla LS does not improve the test accuracy while employing the alternating mechanism yields improvement. CoOp with both CSL and ISL shows higher test accuracy than CoOp and converges earlier. Building upon this, CoOp with ATLaS-CSL and ATLaS-ISL outperforms CoOp with CSL and ISL, respectively. This improved performance suggests that the consistency of labels used in the pre-training and fine-tuning phases can lead to better generalization.
@@
Table 4: Ablation study for the designed soft labels and alternating mechanism. ‘+ $\mathbf{y}$’ denotes co-optimization with one-hot label $\mathbf{y}$, e.g., ‘CSL + $\mathbf{y}$’ is to jointly optimize the prompts using two cross-entropy losses constructed by CSL and the one-hot label $\mathbf{y}$ in each iteration.
Due to page limit, the sensitivity analysis of hyperparameters (i.e., $\theta$, $\tau_{c}$, $\alpha$, and $K$) and more analysis of ATLaS can be found in Appendix D.
5 Conclusion

### v2

@@ IV-A Experimental Setup
Following CoOp [16] and CoCoOp[33], experiments under four settings, including cross-dataset generalization, base-to-new generalization, and few-shot classification, are conducted on 11 image classification datasets. Those datasets include ImageNet [42] and Caltech101 [43] for generic object classification, OxfordPets [44], StanfordCars [45], Flowers102 [46], Food101 [47], and FGVCAircraft [48] for fine-grained visual categorization, SUN397 [49] for scene recognition, DTD [50] for texture classification, EuroSAT [51] for satellite image classification, and UCF101 [52] for action recognition. Besides, following [33], experiments under the domain generalization setting are conducted on the ImageNet dataset and its variants, including ImageNetV2 [53], ImageNet-Sketch [54], ImageNet-A [55], and ImageNet-R [56].
Baselines
@@
We evaluate the effectiveness of the proposed ATLaS method by incorporating it into a broad spectrum of baseline methods, including both textual prompt tuning methods (e.g., CoOp [16], Conditional Context Optimization (CoCoOp) [33], and Decoupled Prompt Tuning (DePT) [35]) and multi-modal prompt tuning methods (e.g., Independent Vision-Language Prompting (IVLP) [57] and Multi-modal Prompt Learning (MaPLe) [34]).
Implementation Details
@@
Our implementation is based on CoOp [16]. For all baseline methods, we maintain the same experimental settings (e.g., training epochs, training schedules, and data augmentation settings) as specified in their original implementations. All experimental results are averaged over three seeds. All experiments are conducted with a ViT-B/16-based CLIP model. The length of prompt vectors is fixed to 4, and they are initialized with the textual template “a photo of a”. For the ATLaS method, $K$, $\theta$ in Eq. (3), $\tau_{c}$ in Eq. (5), and $\alpha$ in Eq. (6) are set to 2, 0.1, 0.05, and 0.1, respectively, by default. More details are provided in Appendix II.
IV-B Experimental Results
To empirically validate the effectiveness of our proposed ATLaS method, we conduct a comprehensive set of experiments across four challenging generalization settings, including cross-dataset generalization, domain generalization, base-to-new generalization, and few-shot classification. We demonstrate that ATLaS and its variants (i.e., ATLaS-CSL that uses the CSL to generate soft labels and ATLaS-ISL that uses ISL to generate soft labels) can be seamlessly integrated into a wide range of prompt tuning methods, including CoOp, CoCoOp, DePT, IVLP, and MaPLe, consistently enhancing their generalization capabilities. The experimental settings for all baselines are identical to those in the original implementations.
IV-B1 Cross-Dataset Generalization
In the cross-dataset generalization setting that evaluates whether prompts learned from a source domain can generalize to unseen target domains, prompts are trained using 16-shot samples from each of the 1000 classes on a single source domain (i.e., ImageNet) and then evaluated on the other 10 unseen target datasets. Results in Table I show that integrating ATLaS and its variants consistently improve the cross-dataset generalization performance over baseline methods. Notably, when integrated with CoOp, ATLaS-CSL achieves the highest average accuracy improvement of 2.10%, showing the powerful benefit of leveraging inter-class textual similarities for domain transfer. The optimal ATLaS variant varies per baseline, highlighting its flexibility in adapting to different prompt tuning architectures.
TABLE I: Performance of various methods under the cross-dataset generalization setting. The number in the parentheses denotes the improvement over the baseline method.
@@ IV-B1 Cross-Dataset Generalization
IV-B2 Domain Generalization
In the domain generalization setting, prompts are evaluated for the ability to generalize to out-of-distribution data. Specifically, prompts are trained using 16-shot samples from each of the 1000 classes on ImageNet, and then evaluated on four different target domains (i.e., ImageNetV2, ImageNet-Sketch, ImageNet-A, and ImageNet-R). Results in Table II show that the prompts learned by the proposed ATLaS method and its variants are more domain-generalizable. Those results further substantiate the effectiveness of the proposed ATLaS method in learning more transferable prompt representations.
TABLE II: Performance of various methods under the domain generalization setting. The number in the parentheses denotes the improvement over the baseline method.
@@ IV-B2 Domain Generalization
CoCoOp 71.02 64.07 48.75 50.63 76.18 59.90
+ ATLaS 70.99 64.21 49.13 51.23 76.99 60.39 (+0.49)
+ ATLaS-CSL 71.05 64.38 49.19 51.41 77.08 60.52 (+0.62)
+ ATLaS-ISL 70.80 64.10 48.90 51.30 77.00 60.33 (+0.43)
DePT 72.79 63.94 45.79 46.51 73.34 57.40
+ ATLaS 72.87 64.03 46.90 47.59 74.48 58.25 (+0.86)
+ ATLaS-CSL 72.83 64.11 46.81 47.63 74.19 58.19 (+0.79)
+ ATLaS-ISL 72.83 64.09 46.85 47.57 74.12 58.16 (+0.76)
IVLP 72.40 65.20 48.80 48.00 76.50 59.63
+ ATLaS 72.60 65.20 49.20 48.70 76.70 59.95 (+0.33)
+ ATLaS-CSL 72.40 65.10 49.20 49.00 76.40 59.93 (+0.30)
+ ATLaS-ISL 71.90 65.30 49.50 48.80 76.60 60.05 (+0.42)
MaPLe 72.01 64.70 48.66 49.21 76.31 59.72
+ ATLaS 71.50 64.53 48.87 50.17 76.97 60.14 (+0.42)
+ ATLaS-CSL 71.47 64.53 48.83 50.23 76.70 60.07 (+0.35)
+ ATLaS-ISL 71.30 64.53 49.30 50.20 77.03 60.27 (+0.55)
IV-B3 Base-to-New Generalization
In the base-to-new generalization setting, all classes are divided into disjoint base and new groups, and the prompts are trained on base classes and evaluated on base (seen) and new (unseen) classes. Table III shows that the proposed ATLaS method consistently enhances the generalization performance of all baselines, specifically achieving better new-class accuracy and harmonic mean accuracy, where detailed results for each dataset are provided in Appendix III. Those results indicate that ATLaS, ATLaS-CSL, and ATLaS-ISL can be seamlessly integrated with existing prompt tuning methods to mitigate overfitting on base classes and enhance generalization to new concepts.
TABLE III: The average performance of five prompt tuning methods with or without ATLaS, ATLaS-CSL, and ATLaS-ISL on the 11 datasets under the base-to-new generalization setting. ‘tp’ denotes textual prompt tuning, and ‘mmp’ denotes multi-modal prompt tuning. ‘H’ denotes the harmonic mean accuracy. The number in the parentheses denotes the improvement over the baseline method.
@@ IV-B3 Base-to-New Generalization
IV-B4 Few-shot Classification
In the few-shot classification setting, prompts are trained using a limited number of labeled samples and evaluated on the remaining samples. We conduct few-shot classification using 1, 2, 4, 8, and 16 samples per class. The indicator function $\delta$ is always set to 1 under the few-shot classification setting. As visualized in Fig. 4, the advantages of the proposed method are pronounced even with very few training examples. ATLaS and its variants consistently outperform the CoOp baseline across all shots in terms of the average performance on the 11 datasets. Those results demonstrate that our method’s ability to enhance generalization is not contingent on large amounts of training data, making it highly valuable for practical, low-resource applications.
Figure 4: Performance of various methods on the 11 datasets under the few-shot classification setting.
IV-C Ablation Studies
@@
In this section, we conduct ablation studies to analyze the effectiveness of the alternating mechanism and the proposed soft labels (i.e., CSL and ISL) in the proposed ATLaS method. The experiments are performed using CoOp [16] as the baseline method under the base-to-new generalization setting.
@@
Table IV shows the average performance of CoOp equipped with different LS strategies on the 11 datasets. Based on the results, we have several observations.
@@
  • CoOp with vanilla LS degrades the performance of CoOp (i.e., -0.45% in the base-class accuracy, -0.91% in the new-class accuracy, -0.75% in the harmonic mean accuracy), suggesting that such a naive combination makes the inconsistency of labels used in the pre-training and fine-tuning phases, weakening generalization performance. In contrast, CoOp with ATLaS, which alternately uses one-hot and soft labels to supervise prompt tuning, outperforms CoOp by 1.96% in the new-class accuracy (i.e., 65.19% vs. 63.23%) and 1.14% in the harmonic mean accuracy (i.e., 72.80% vs. 71.66%), demonstrating that such alternating fine-tuning can enhance prompt generalization.
@@
  • Supervising prompt tuning with CSL and ISL as soft labels improves the new-class accuracy (i.e., +2.5% for CSL, +10.41% for ISL) and harmonic mean accuracy (i.e., +0.64% for CSL, 3.58% for ISL) at the expense of decreasing the base-class accuracy (i.e., -2.36% for CSL, -5.79% for ISL). This occurs because both CSL and ISL leverage the inherent properties of CLIP, allowing them to enhance prompt generalization while preserving the inherent knowledge of CLIP. Building upon this foundation, the alternating mechanism further improves their performance, that is, compared with CoOp, ATLaS-CSL and ATLaS-ISL improve the new-class accuracy by +4.82% and +10.39% respectively, and harmonic mean accuracy by +2.73% and +5.13%. Those results indicate that both the alternating mechanism and soft labels can enhance prompt generalization.
@@
  • Alternating optimization outperforms joint optimization when using both one-hot labels and soft labels to train prompts. That is, CoOp with ATLaS outperforms CoOp with ‘vanilla LS + y’ in terms of the harmonic mean accuracy (i.e., 72.80% v.s. 71.68%), further demonstrating the effectiveness of the proposed alternating mechanism.
TABLE IV: Ablation study for the designed soft labels and alternating mechanism. The average performance of 11 datasets is reported. Prompts are trained with 16 instances per base class. ‘+ $\mathbf{y}$’ denotes co-optimization with one-hot label $\mathbf{y}$, e.g., ‘CSL + $\mathbf{y}$’ is to jointly optimize the prompts using two cross-entropy losses constructed by CSL and the one-hot label $\mathbf{y}$ in each iteration. The number in the parentheses denotes the improvement or degradation over the baseline method.
@@
Fig. 5 shows the test accuracy of CoOp and CoOp with vanilla LS, ATLaS, CSL, ATLaS-CSL, ISL, and ATLaS-ISL on new classes of the SUN397 dataset over the training process. The test accuracy of CoOp first increases and then drops sharply. Integrating CoOp with vanilla LS does not improve the test accuracy while employing the alternating mechanism yields improvement. CoOp with both CSL and ISL shows higher test accuracy than CoOp and converges earlier. Building upon this, CoOp with ATLaS-CSL and ATLaS-ISL outperforms CoOp with CSL and ISL, respectively. This improved performance suggests that the consistency of labels used in the pre-training and fine-tuning phases can lead to better generalization.
IV-D Sensitivity Analysis
In this section, we investigate the impact of four hyper-parameters (i.e., $\theta$, $\tau_{c}$, $\alpha$, and $K$) on the average performance across 11 datasets under the base-to-new generalization setting and the results are shown in Fig. 6.
Figure 6: The average performance of ATLaS (a), ATLaS-CSL (b), ATLaS-ISL (c, d) on all datasets with respect to hyper-parameters (i.e., $\theta$, $\tau_{c}$, $\alpha$, and $K$) under the base-to-new generalization setting.
Impact of the smoothing parameter $\theta$. The smoothing parameter $\theta$ in Eq. (3) balances one-hot labels against the uniform distribution. Fig. 6a shows the average performance of CoOp with ATLaS across the 11 datasets for $\theta$ values in {0.1,…,0.6}. As $\theta$ increases, soft labels exhibit reduced confidence in target classes and increased attention to non-target classes. This leads to a gradual decline in the base-class accuracy and improvement in the new-class accuracy, while the harmonic mean accuracy remains relatively stable.
Impact of the temperature parameter $\tau_{c}$. The temperature parameter $\tau_{c}$ in Eq. (5) controls the smoothness of the similarity distribution across inter-class textual embeddings. A lower $\tau_{c}$ sharpens the distribution, enhancing sensitivity to inter-class distinctions, while a higher value smooths the distribution. Fig. 6b shows the average performance of CoOp with ATLaS-CSL across 11 datasets for $1/\tau_{c}$ values in {10, 20,…, 100}. It can be seen that as $\tau_{c}$ decreases, the base-class accuracy initially increases and subsequently stabilizes, while both the new-class accuracy and the harmonic mean accuracy exhibit an initial increase followed by a decline.
Impact of the rectification coefficient $\alpha$. The rectification coefficient $\alpha$ in Eq. (6) controls how significantly the one-hot label modifies the soft label. Higher $\alpha$ values intensify the rectification applied to the prediction probabilities of CLIP. Fig. 6c shows the average performance of CoOp with ATLaS-ISL across 11 datasets for $\theta$ values in {0.1, 0.2,…, 1}. As $\alpha$ increases, the base-class accuracy gradually improves, while the new-class accuracy initially increases, then declines gently. The harmonic mean accuracy remains generally stable across this range.
Impact of the alternating period $K$. The alternating period $K$ regulates the ratio between epochs using one-hot labels and soft labels for prompt tuning. Fig. 6d shows the average performance of CoOp with ATLaS-ISL across 11 datasets for $K$ in {1, 2,…, 5}. Note that when $K=1$, the alternating training mechanism is disabled, and only instance-wise soft labels are used for training. As $K$ increases, the base-class accuracy gradually improves, while the new-class accuracy decreases, which aligns with our expectation since a larger $K$ will increase the proportion of one-hot labels used during training. Notably, the harmonic mean accuracy with the alternating mechanism (i.e., $K>1$) consistently outperforms that without the alternating mechanism (i.e., $K=1$), validating the effectiveness of the proposed alternating mechanism.
IV-E Combination of CSL and ISL
We explored CSL and ISL mixing ($w\times\mathrm{CSL}+\left(1-w\right)\times\mathrm{ISL}$). As shown in Fig. 7, on ImageNet, ATLaS achieves the best performance at $w=0.3$, surpassing individual CSL and ISL. However, we found that this mixing approach does not outperform standalone ISL or CSL across all scenarios.
Figure 7: ATLaS with mixed ISL and CSL on the ImageNet dataset under the base-to-new generalization setting.
IV-F Further Analysis of the Alternating Period $K$
As mentioned in Sec. III-C, vanilla LS uses a uniform distribution to generate low-quality soft labels ($\kappa>1$). In such case, the alternating mechanism can improve the performance of vanilla LS. Figure 8 shows that when the alternating training mechanism is used (i.e., $K>1$), ATLaS maintains advantages over vanilla LS.
However, when high-quality soft labels (e.g., ISL) are used, $\kappa$ can be less than 1. Under this condition, Theorem 1 indicates that non-alternating LS achieves a tighter convergence bound than ATLaS. Therefore, Figure 6d also aligns with our theoretical analysis, where $K=1$ (using ISL without alternating training) achieves better new-class accuracy compared to other $K$ using the alternating mechanism. Nevertheless, the alternating mechanism can enhance the overall performance. Figure 6d shows that the highest H score, which reflects balanced effectiveness on both Base and New classes, is achieved at $K=2$. This occurs because ATLaS-ISL improves base-class accuracy through one-hot label training while preventing overfitting to base classes.
Figure 8: The average performance of ATLaS on all datasets with different $K$ under the base-to-new generalization setting
IV-G Computational Cost
[…]
TABLE V: Computational cost of the proposed ATLaS, where the model is trained on the base classes of ImageNet with 16-shot examples per class.
Method Train. time Infer.time H (Avg.)
CoOp 1746s 2.33ms 71.66
+ATLaS +0s +0ms 72.30
+ATLaS-CSL +1.96s +0ms 74.39
+ATLaS-ISL +2.47s +0ms 76.79
V Conclusion